OpenStudy (anonymous):

can someone please explain how to do this problem - step by step: P=2L+2W solve for W

OpenStudy (bahrom7893):

P = 2L + 2W Subtract 2L from both sides: P - 2L = 2L + 2W - 2L P - 2L = 2L - 2L + 2W P - 2L = 2W

OpenStudy (bahrom7893):

Divide both sides by 2: (P - 2L)/2 = 2W/2

OpenStudy (bahrom7893):

(P/2) - L = W, or W = (P/2) - L

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

I found you